# Evaluating Fairness of Voting Systems: Simulating Violations of Arrow’s Conditions

### v3

@@ Abstract
This paper builds upon the work of Dougherty and Heckelman (2020) by computing violation frequencies for elections with a greater number of alternatives. These results determine which of the 13 voting systems, or preference aggregation rules, is the fairest based on their probabilistic likelihood of violating Arrow’s social choice criteria. The voting systems considered are: Plurality, Borda, Dowdall, Top Two, Instant Runoff, Coombs, Baldwin, Copeland, Anti-Plurality, Nanson, Ranked Pairs, Pairwise Majority, and Minimax. Elections with up to 10,000 voters and between three and six alternatives are simulated using both Impartial Culture and Impartial Anonymous Culture. The results of these simulations produce new IIA violation frequencies for each method and show that Pairwise Majority is the least likely to jointly violate Arrow’s criteria. As the number of alternatives increases, the joint-violation frequencies increase for each voting method. For all systems except Pairwise Majority, the joint-violation frequencies in elections with at least 30 voters and four alternatives are greater than 98%.
@@ 1 Introduction
Social choice theory studies how individual preferences can be aggregated into collective decisions, often through voting systems. A central challenge in this field is determining whether any such system can satisfy desirable fairness criteria. Kenneth Arrow’s Impossibility Theorem (1963) famously demonstrated that no rank-order voting system can satisfy a set of intuitive fairness conditions - namely, Non-Dictatorship, Unrestricted Domain, Transitivity, Unanimity, and Independence of Irrelevant Alternatives (IIA) - when there are at least three alternatives and two or more voters. While the theorem is negative in nature, it has motivated decades of research focused on evaluating which voting systems are least likely to violate these criteria in practice, especially under probabilistic models of voter behavior (e.g., [22]; [20]).
@@ 2 Voting Systems
As our work extends the results of Dougherty and Heckelman (2020) to more alternatives, we consider the same set of voting systems defined equivalently below. Note that Plurality, Minimax, Top Two, Instant Runoff, Anti-Plurality, and Nanson are said to be part of Group 1 due to their similar numerical results.
@@
Alternatives are ranked in descending order based on their first-place vote tally.
@@
Alternatives receive zero points for every last place vote and 1 point for all other place votes. Alternatives are ranked in descending order based on their point total.
@@
Let the election have $n$ alternatives. For each 1st place vote, alternatives receive $n$ points, for each 2nd place vote they receive n1n-1 points, and so on until for an $n$th place vote, alternatives receive one point. Alternatives are ranked in descending order based on their point total.
@@
Let the election have $n$ alternatives. For each first place vote, alternatives receive 1 point, for a second place vote they receive $\frac{1}{2}$ point, for a third place vote they receive $\frac{1}{3}$ point and so on until for an $n$th place vote, alternatives receive $\frac{1}{n}$ point. Alternatives are ranked in descending order based on their point total.
@@
All alternatives are compared head to head pairwise. For two alternatives $A$ and $B$, if $A$ wins the head to head then $A\succ B$. If $B$ wins the head to head then $B\succ A$. If they tie then $A\sim B$. Rank alternatives based on the pairwise comparisons.
@@
All alternatives are compared head to head pairwise. For each comparison, the alternative that wins head to head receives 1 point. If the two alternatives tie they receive $\frac{1}{2}$ point each. Rank alternatives in descending order of total points.
@@ 2.1 Arrow’s Criteria of Fairness
Following [20, 25, 31], below is a list of Arrow’s voting criteria used in this paper. Let $C=\{1,2,\dots,c\}$ be a finite set of alternatives and $V=\{1,2,\dots,v\}$ be a finite set of voters.
@@ 3 Arrow’s Theorem
For a formal proof of the above results, see [32].
@@ 4 Simulations
For Top Two, there is a possibility of a tie in the first round, where more than two alternatives have an equal number of votes. In this case, ties are broken randomly. In all other methods, if alternatives are tied, they have the same rank in the societal preference order.
Two preference distributions, Impartial Culture and Impartial Anonymous Culture [21] are considered. Impartial Culture (IC) assumes each voter is equally likely to select any of the possible preference orders. Impartial Anonymous Culture (IAC) assumes each preference schedule is equally likely to occur. All pseudo-random number generation was done using the Mersenne Twister engine [24] in C++ random library.
@@ 5 Results and Analysis
The following sections detail the likelihood of voting systems violating Arrow’s social choice criterion. Note that all voting methods considered satisfy Non-Dictatorship and Unrestricted Domain and as such these criteria are excluded from the analysis.
@@ 5.1 Independence of Irrelevant Alternatives (IIA)
Under IAC, Ranked Pairs has the lowest violation frequencies. For elections with 10,000 voters and 3 alternatives, Ranked Pairs violates IIA at a frequency of 75.87% for 10000 voters. For the same election, Coombs, Borda, and Baldwin all violate IIA at frequencies between 83% and 86%.
@@ 5.2 Pareto Unanimity (U)
The results for Instant Runoff, Plurality, and Top Two are identical. Data for all Group 1 systems is identical for three alternatives, but differ for four or more alternatives as noted in literature [20]. For larger number of alternatives, Nanson and Minimax perform the best, with less than 30% violations with six alternatives.
@@ 5.4 Joint Violation of Arrow’s Criteria
Simulations were run to find the percentage of times a voting system violates U, IIA, and TR. The other criterion (ND, UD) were not simulated. The joint-violation (JV) frequencies, defined as the percentage of times systems violated any one criteria, serve as an indication of fairness of a voting system.
@@ 6 Conclusion
This study contributes to the enduring challenge in social choice theory of identifying voting systems that perform best with respect to Arrow’s axiomatic criteria. While Arrow’s impossibility theorem demonstrates that no voting system can simultaneously satisfy all fairness conditions under general preference domains, our results provide an empirical assessment of how often different systems fail in practice. Extending prior work by Dougherty and Heckelman (2020), our simulations incorporate larger electorates and a broader range of alternatives, offering a more comprehensive picture of violation frequencies across commonly used preference aggregation rules.
Our central finding - that Pairwise Majority exhibits the lowest joint-violation frequency among all tested systems -reinforces prior theoretical observations (Gehrlein, 1994; Niemi & Weisberg, 1968) and suggests that even within the constraints of Arrow’s theorem, some systems may better approximate normative ideals in practical settings. Importantly, we demonstrate that the likelihood of violating Arrow’s conditions increases as the number of alternatives increases, a trend consistent with both theoretical expectations and empirical findings (Nurmi, 1983; Börgers, 2010). These insights are especially relevant in contemporary democratic societies where elections increasingly involve crowded candidate fields and diverse voter preferences.
@@ A.2 Unanimity
Table 14: Percentage U Violations Plurality, Top Two, Instant Runoff

### v4

@@ Abstract
This paper builds upon the work of Dougherty and Heckelman (2020) by determing the frequency that 13 voting systems violate Arrow’s social choice criteria with up to six alternatives. These results determine which of the 13 voting systems, is the fairest based on their probabilistic likelihood of violating Arrow’s social choice criteria. The voting systems considered are: Plurality, Borda, Dowdall, Top Two, Hare, Coombs, Baldwin, Copeland, Anti-Plurality, Nanson, Ranked Pairs, Pairwise Majority, and Minimax. Elections with up to 10,000 voters and between three and six alternatives are simulated using both Impartial Culture and Impartial Anonymous Culture. These simulations show that Pairwise Majority is the least likely to jointly violate Arrow’s criteria. As the number of alternatives increases, the joint-violation frequencies increase for each voting method. For all systems except Pairwise Majority, the joint-violation frequencies in elections with at least 30 voters and four alternatives are greater than 98%.
@@ 1 Introduction
Social choice theory studies how individual preferences can be aggregated into collective decisions, often through voting systems. A central challenge in this field is determining whether any such system can satisfy desirable fairness criteria. Kenneth Arrow’s Impossibility Theorem (1963) famously demonstrated that no rank-order voting system can satisfy a set of intuitive fairness conditions - namely, Non-Dictatorship, Unrestricted Domain, Transitivity, Unanimity, and Independence of Irrelevant Alternatives (IIA) - when there are at least three alternatives and two or more voters. While the theorem is negative in nature, it has motivated decades of research focused on evaluating which voting systems are least likely to violate these, and other, criteria in practice, especially under probabilistic models of voter behavior (e.g., [24]; [22]).
@@ 2 Voting Systems
Let $C=\{1,2,\dots,m\}$ be a finite set of alternatives and $V=\{1,2,\dots,v\}$ be a finite set of voters.
We consider the same set of voting systems as defined by Dougherty and Heckelman (2020). Following their lead we classify Plurality, Minimax, Top Two, Hare, Anti-Plurality, and Nanson as Group 1 because those voting systems always violate IIA and can violate Pareto - two of Arrow’s conditions defined below.
@@
Alternatives are ranked in descending order based on their first-place vote tally. Alternatives with the same number of first place votes are given the same rank.
@@
Alternatives receive zero points for every last place vote and 1 point for all other place votes. Alternatives are ranked in descending order based on their point total. Alternatives
@@
For each 1st place vote, alternatives receive $m$ points, for each 2nd place vote they receive m1m-1 points, and so on until for an $m$th place vote, alternatives receive one point. Alternatives are ranked in descending order based on their point total.
@@
For each first place vote, alternatives receive 1 point, for a second place vote they receive $\frac{1}{2}$ point, for a third place vote they receive $\frac{1}{3}$ point and so on until for an $m$th place vote, alternatives receive $\frac{1}{m}$ point. Alternatives are ranked in descending order based on their point total.
@@
Let $A,B\in C$ be two distinct alternatives. The head-to-head comparison between $A$ and $B$ is determined by how many voters rank $A$ above $B$ compared to how many rank $B$ above $A$. If $A$ defeats $B$ when compared head-to-head, then $A\succ B$. If $B$ defeats $A$ when compared head-to-head then $B\succ A$. If they tie when compared head-to-head, then $A\sim B$. Rank alternatives based on the pairwise comparisons.
@@
All alternatives are compared head-to-head pairwise. For each comparison, the alternative that wins head-to-head receives 1 point. If the two alternatives tie they receive $\frac{1}{2}$ point each. Rank alternatives in descending order of total points.
@@
2.1 Breaking Ties
In Plurality, Anti-Plurality, Borda Count, Dowdall, Copeland, and Minimax, alternatives with the same number of points or votes are assigned the same rank. In round based systems, such as Hare, Coombs, Baldwin, and Nanson, alternatives eliminated in the same round are assigned the same rank. In Pairwise Majority, if $A$ and $B$ are tied in the head-to-head comparison then $A\sim B$ in the final societal ordering. For Ranked Pairs, if two or more pairs share the largest vote margin, one pair is randomly selected. For Top Two, a tie may occur in the first round if more than two alternatives receive an equal number of votes. In this case, ties are broken randomly.
@@ 2.2 Arrow’s Criteria of Fairness
Following [22, 27, 33], below is a list of Arrow’s voting criteria used in this paper.
@@ 3 Arrow’s Theorem
For a formal proof of most of the above results, see [34].
@@ 4 Simulations
Two preference distributions, Impartial Culture and Impartial Anonymous Culture [23] are considered. Impartial Culture (IC) assumes each voter is equally likely to select any of the possible strict preference orders. Impartial Anonymous Culture (IAC) assumes each preference schedule is equally likely to occur. All pseudo-random number generation was done using the Mersenne Twister engine [26] in C++ random library.
@@ 5 Results and Analysis
The following sections detail the likelihood of voting systems violating Arrow’s social choice criterion. As preference schedules are drawn from IC and IAC uniform distributions, UD is satisfied by assumption. The analysis for the remaining criteria is given below. ND is excluded as all 13 voting systems satisfy ND.
@@ 5.1 Independence of Irrelevant Alternatives (IIA)
Among the voting systems compared, Ranked Pairs violates IIA least frequently. For elections with 10,000 voters and 3 alternatives, Ranked Pairs violates IIA at a frequency of 75.87% for 10000 voters. For the same election, Coombs, Borda, and Baldwin all violate IIA at frequencies between 83% and 86%.
@@ 5.2 Pareto Unanimity (U)
The results for Hare, Plurality, and Top Two are identical. Data for all Group 1 systems is identical for three alternatives, but differ for four or more alternatives as noted in literature [22]. For larger number of alternatives, Nanson and Minimax perform the best among the group, with less than 30% violations with six alternatives.
@@ 5.4 Joint Violation of Arrow’s Criteria
Simulations were run to find the percentage of times a voting system violates U, IIA, and TR. The joint-violation (JV) frequencies, defined as the percentage of times a system violated at least one criterion, serve as an indication of fairness of a voting system.
@@ 6 Conclusion
This study contributes to the enduring challenge in social choice theory of identifying voting systems that perform best with respect to Arrow’s axiomatic criteria. While Arrow’s impossibility theorem establishes that no voting system can simultaneously satisfy all fairness conditions under general preference domains, our results provide an empirical assessment of how often different systems fail in practice. Extending prior work by Dougherty and Heckelman (2020), our simulations consider larger electorates and a wider range of alternatives, offering a more comprehensive picture of violation frequencies across commonly used preference aggregation rules.
Our central finding - that Pairwise Majority exhibits the lowest joint-violation frequency among all tested systems - reinforces prior theoretical observations (Gehrlein, 1994; Niemi & Weisberg, 1968) and suggests that even within the constraints of Arrow’s theorem, some systems may better approximate normative ideals in practical settings. Importantly, we demonstrate that the likelihood of IIA violations rises sharply with the number of alternatives, a trend consistent with both theoretical expectations and empirical findings (Nurmi, 1983; Börgers, 2010). Since almost every voting rule is susceptible to IIA violations, this result highlights their ubiquity and underscores that systems must often be judged on other normative grounds unless pairwise majority rule is employed.
These insights are especially relevant in contemporary democratic societies where elections increasingly involve crowded fields and diverse voter preferences. At the same time, we caution that our conclusions are based on simulations under the uniform distribution of preferences; different distributions, particularly skewed ones, may yield different violation probabilities.
@@ A.2 Unanimity
Table 14: Percentage U Violations Plurality, Top Two, Hare